# Skill-informed Data-driven Haptic Nudges for High-dimensional Human Motor Learning

### 13 Mar 2026 v1

@@ Abstract
In this work, we propose a data-driven skill-informed framework to design optimal haptic nudge feedback for high-dimensional novel motor learning tasks. We first model the stochastic dynamics of human motor learning using an Input-Output Hidden Markov Model (IOHMM), which explicitly decouples latent skill evolution from observable kinematic emissions. Leveraging this predictive model, we formulate the haptic nudge feedback design problem as a Partially Observable Markov Decision Process (POMDP). This allows us to derive an optimal nudging policy that minimizes long-term performance cost, implicitly guiding the learner toward robust regions of the skill space. We validated our approach through a human-subject study ($N=30$) using a high-dimensional hand-exoskeleton task. Results demonstrate that participants trained with the POMDP-derived policy exhibited significantly accelerated task performance compared to groups receiving heuristic-based feedback or no feedback. Furthermore, synergy analysis revealed that the POMDP group discovered efficient low-dimensional motor representations more rapidly.
@@ 1 Introduction
The integration of robotic systems into human spaces presents a transformative opportunity to improve humans’ quality of life through automated and personalized physical assistance during skill acquisition and motor rehabilitation. Haptic feedback has proven to be a powerful modality to guide human motor motions, thus improving task performance and skill learning without fostering dependence on robotic assistance [8, williams2014motor]. However, learning to perform high-dimensional de-novo (novel) tasks, like using a tool or recovering motor function post neurological injury, is much harder than simple low-dimensional reaching movements [16]. Consequently, designing haptic assistance for such complex tasks is an open challenge since determining the optimal/perfect movement for such tasks is non-trivial; de-novo learning involves discovering a movement strategy, rather than simply reducing task error.
@@
Haptic guidance designs for physical human robot interaction regimes can be broadly categorized into two approaches: using expert-guidance or based on observed task performance metrics. The first category involves field experts and professionals designing haptic guidance systems informed by their intuition, experience, and knowledge [turolla2013haptic, lin2016development, esmatloo2020fingertip]. This “top-down” approach is not suitable for high degrees of freedom (DoF) tasks where the solution space is vast, owing to redundancy in motor space, and finding optimal motions to instruct the right motor coordination strategies is non-trivial [11]. The second category exploits directly observable task performance metrics like task errors, movement time, smoothness, or task success [basalp2021haptic, sigrist2013augmented, sullivan2021haptic, rowe2017robotic]. While this reactive strategy is intuitive and easy to design, it can often be misleading as performance is a noisy measure of true skill, especially for high DoF tasks. The assistance design might overfit to the user, helping them perform well temporarily (low errors) by using a bad, compensatory habit (low skill). The simplified task performance might not capture important features of the underlying latent skill state, thus leading learners toward suboptimal solutions. This prevents mastery of high-dimensional skill and ultimately hinders long-term skill development.
@@
To address these limitations, we propose haptic assistance designs informed by latent skill rather than task performance corrections. Central to our approach is the recognition that motor learning is a latent process, where observable kinematic performance is a noisy emission of the underlying evolving cognitive control strategy [14]. Drawing on cognitive diagnosis and human skill modeling literature [17, 14, 18, 15], we formalize motor learning as a progression through discrete latent states, paralleling the cognitive, associative, and autonomous phases proposed by Fitts and Posner [1]. While standard Hidden Markov Models (HMM) have successfully tracked such transitions in cognitive tasks [14] and neural motor planning [4, 5], effective robotic nudge feedback requires understanding how external interventions influence these transitions. Thus, we employ a probabilistic predictive model of human motor learning using an Input-Output Hidden Markov Model (IOHMM). This structure allows us to causally link specific haptic nudges (inputs) to probabilities of skill improvement, effectively separating execution noise from true latent capability.
@@
Several mechanistic models have been proposed in the literature to explain the doubly stochastic human motor learning process [13, 10, 3]. One salient feature of these models is the Markovian evolution of internal/latent states as a function of certain inputs to the model; these models can be discretized to obtain the discrete-time update rule for these latent states. Further binning the continuous latent state space into discrete skill states gives us the standard IOHMM structure. This probabilistic framework allows us to treat the haptic tutoring as a sequential decision problem making under uncertainty. Leveraging the IOHMM, we formulate a Partially Observable Markov Decision Process (POMDP) to derive an optimal nudging policy. Similar to prior work that utilized POMDPs to infer latent user fatigue during rehabilitation [2], our formulation models the evolution of latent motor skill over the duration of the motor task being performed by the learner. The objective is to prescribe a sequence of haptic feedback nudges that minimizes long-term performance cost, implicitly guiding the learner toward effective regions of the latent skill space. This enables a principled, data-driven approach, going beyond performance or expert-designed heuristics, for personalizing haptic feedback design in novel skill acquisition paradigms.
@@
Building on the preliminary findings of our POMDP-based optimal nudge feedback [kamboj2026toward], this work investigates the potential of the skill-informed haptic feedback design in accelerating skill acquisition and task performance for high-dimensional, de-novo motor tasks. We distinguish this study from previous work through the following advancements: (i) we introduce another experiment group training on a heuristics-based nudging policy to benchmark performance against the POMDP group, (ii) we increase the sample size to better delineate group effects, (iii) we employ additional analysis to evaluate how effectively different group participants are discovering the task space manifold, (iv) we evaluate the learning dynamics to analyze the learning trajectories and their convergence for the experiment groups, (v) we characterize the participants’ task performance using the reaching error metric, and (vi) we critically discuss the data-driven motor learning behavioral model and provide interpretability to the estimated model parameters. In summary, our contributions are threefold as follows.
@@
    We introduce a generalizable IOHMM framework that captures the dynamics of high-dimensional motor learning processes as a function of haptic feedback and task difficulty, explicitly decoupling unobservable skill from noisy performance.
@@
    We validate our approach through a human-subject study, demonstrating that our POMDP-derived policy significantly expedites motor skill learning and enhances task performance outcomes compared to both no-feedback baselines and heuristic-based feedback strategies.
@@
Figure 1: Target Capture Game: (a) shows a participant playing our target capture game with the SenseGlove DK1 exoskeleton strapped to their right hand. (b) and (c) show the cursor trajectories during the first and eighth blocks of the game with the red dots representing the target locations. The trajectories get straighter as the participant learns to control the cursor by the end of the experiment session. (d) shows how the performance metrics, Reaching Error and Straightness of Trajectory, are computed.
@@ 2 Background
This section details the experimental framework and the performance metrics used to evaluate human performance during the acquisition of a novel, high-dimensional motor skill. Following the experimental description, we outline the stochastic modeling approach, specifically a Markov chain framework employed to characterize how control interventions influence the user’s latent skill state and subsequently modulate observable task performance.
@@ 2.1 Experimental Setup
Our study is grounded in the established motor learning principle that the neural mechanisms driving de-novo skill acquisition are analogous to those involved in motor recovery following neurological injuries, such as stroke [6]. To operationalize this, we adopt a Body-Machine Interface (BoMI) paradigm similar to those described in [9, 11]. Participants wore a SenseGlove DK1 (manufactured by SenseGlove, a Netherlands-based company), a non-invasive hand exoskeleton capable of tracking $20$ distinct degrees of freedom (DoF) across the hand joints. Additionally, the device can deliver independent vibrotactile haptic feedback to each of the five fingertips. This high-dimensional joint space is then projected into a lower-dimensional space, represented by a cursor on a screen. This projection is governed by the linear mapping
@@
where $\dot{\boldsymbol{x}}\in\mathbb{R}^{n}$ represents the velocity of the screen cursor, $\boldsymbol{u}=\dot{\boldsymbol{q}}\in\mathbb{R}^{m}$ denotes the finger joint velocities derived from the joint angles $\boldsymbol{q}$, and $C\in\mathbb{R}^{n\times m}$ serves as the BoMI mapping matrix. In the experimental task, referred to as the target capture game [^1] A video footage of the target capture task is available at https://youtu.be/WIDRwqpE9Sk (see Fig. 1), participants are instructed to coordinate their finger joint movements to navigate the cursor toward sequentially appearing targets. The targets are prescribed uniformly randomly from a set of $4$ targets $\mathcal{T}=\{(0.5,4.5),(2.5,2.5),(2.5,0.5),(4.5,4.5)\})$ units of the game window grid. Because the mapping from the $20$-dimensional hand space to the $2$-dimensional cursor space is initially unknown to the user, the task enforces a high-dimensional learning process. Success requires the participant to discover and internalize specific coordinated hand motions consistent with the mapping matrix $C$. The experimental protocol is the same as that adopted in [kamboj2026toward], and we omit it here for brevity.
@@ 2.2 Performance Metrics
To quanepsy motor performance, we utilize two primary metrics [3]: Reaching Error (RE) and Straightness of Trajectory (SoT), as shown in Fig. 1(d).
  • RE is defined as the Euclidean distance between the cursor and the target center at the conclusion of a trial. The trial endpoint is determined by the stability condition described above, or a time-out of $2$ seconds, whichever occurs first.
@@ 2.3 IOHMM and POMDP Framework
To model the stochastic dynamics of the user’s internal skill-state, we employ an Input-Output Hidden Markov Model (IOHMM) [bengio1996input]. This probabilistic graphical model describes the evolution of a discrete latent state sequence $\{s_{k}\in\mathcal{S}\}_{k\in\mathbb{N}}$, where the transition to a future state depends on both the current state $s_{k-1}$ and an exogenous control input or action $a_{k}\in\mathcal{A}$. Crucially, the true system state is inaccessible for direct measurement; it is hidden. Instead, inference is performed via observable emissions $\{o_{k}\in\Omega\}_{k\in\mathbb{N}}$. These observations are generated stochastically from the conditional probability distribution $\mathbb{P}(o|s,a)$, representing the likelihood of observing output $o$ given the system is in state $s$ and receives input $a$. We assume the state space $\mathcal{S}$ and action space $\mathcal{A}$ are finite sets.
The primary objective in training an IOHMM is system idenepsication: estimating the underlying parameters given a historical dataset of input-output pairs $\{a_{k},o_{k}\}_{k\in\mathbb{N}}$. Specifically, we seek to learn the state transition dynamics $T(s,a,s^{\prime})=\mathbb{P}(s^{\prime}|s,a)$, the observation probability matrix $O(s,a,o)=\mathbb{P}(o|s,a)$, and the initial state distribution. This parameter estimation is typically achieved via the extended Baum-Welch algorithm [bengio1996input], which adapts the standard Expectation-Maximization (EM) algorithm to account for the input-dependent structure of the IOHMM.
To transition from system modeling to active control, we extend the IOHMM into a Partially Observable Markov Decision Process (POMDP). The POMDP framework introduces an agent that optimizes the selection of inputs based on a defined objective. Formally, the POMDP is characterized by the tuple $\langle\mathcal{S},\mathcal{A},T,R,\Omega,O\rangle$. In this formulation, $\mathcal{S},\mathcal{A},T,\Omega,$ and $O$ retain their definitions from the IOHMM, while $R(s,a)$ represents the reward function. This function quanepsies the agent’s performance goals, assigning positive scalar values to beneficial state-action pairs and penalties (costs) to undesirable outcomes, thereby guiding the optimal control policy.

### 13 Apr 2026 v2

@@ Abstract
In this work, we propose a data-driven framework to design optimal haptic nudge feedback leveraging the learner’s estimated skill to address the challenge of learning a novel motor task in a high-dimensional, redundant motor space. A nudge is a series of vibrotactile feedback delivered to the learner to encourage motor movements that aid in task completion. We first model the stochastic dynamics of human motor learning under haptic nudges using an Input-Output Hidden Markov Model (IOHMM), which explicitly decouples latent skill evolution from observable performance measures. Leveraging this predictive model, we formulate the haptic nudge feedback design problem as a Partially Observable Markov Decision Process (POMDP). This allows us to derive an optimal nudging policy that minimizes long-term performance cost and implicitly guides the learner toward superior skill states. We validate our approach through a human participant study ($N=30$) involving a high-dimensional motor task rendered through a hand exoskeleton. Results demonstrate that participants trained with the POMDP-derived policy exhibit significantly accelerated movement efficiency and endpoint accuracy compared to groups receiving heuristic-based feedback or no feedback. Furthermore, synergy analysis reveals that the POMDP group discovers efficient low-dimensional motor representations more rapidly.
@@ 1 Introduction
The integration of robotic systems into human environments presents a transformative opportunity to support skill acquisition and motor learning through automated and personalized physical assistance. However, facilitating learning in high-dimensional de-novo (novel) tasks, such as manipulating a complex tool or recovering motor function post neurological injury, poses a formidable computational and cognitive challenge [22]. De-novo learning necessitates the discovery of novel movement strategies, which fundamentally distinguishes it from the simple reduction of kinematic task errors observed in low-dimensional reaching tasks [5, 10]. Furthermore, the human motor system exhibits massive kinematic redundancy, meaning the solution space for any high degrees of freedom (DoF) task is vast [17]. Consequently, learners must navigate this highly complex manifold to identify optimal motor coordination strategies. Designing automated guidance for this learning process is an open challenge since determining the optimal movement for such tasks is non-trivial.
@@
To assist in this complex skill acquisition process, haptic feedback has proven to be a powerful interventional modality that guides human movements without fostering dependence on robotic assistance [14, williams2014motor]. Haptic guidance designs for physical human robot interaction regimes can be broadly categorized into two approaches: (i) expert-in-the-loop, and (ii) task-performance-based automated approaches. The first category involves field experts and professionals designing haptic guidance systems informed by their intuition, experience, and knowledge [turolla2013haptic, lin2016development, esmatloo2020fingertip]. While effective for constrained movements, this “top-down” approach scales poorly to high-DoF tasks where finding optimal motions to instruct precise motor coordination strategies is overwhelmingly complex due to the aforementioned redundancy. The second category exploits directly observable task performance metrics like task errors, movement time, smoothness, or task success to design haptic robotic feedback [basalp2021haptic, sigrist2013augmented, sullivan2021haptic, rowe2017robotic]. Although this reactive haptic feedback strategy is intuitive and easier to automate, it can often be misleading as performance is a noisy measure of true skill, especially for high DoF tasks. This assistance might overfit to the user, helping them perform well temporarily (low errors) by using a bad, compensatory habit (low skill). The simplistic task performance might not capture important features of the underlying latent skill state, thus leading learners toward suboptimal solutions. This prevents mastery of high-dimensional skill and ultimately hinders long-term skill development.
To address these limitations, we propose automated haptic nudge assistance designs informed by latent skill rather than task performance corrections. In our work, the robotic nudge feedback is a series of two vibrations delivered to a finger during the early phase of a trial, suggesting that the participant should move that finger more to accomplish the task. We specifically employ discrete vibrotactile nudges over continuous force feedback to maintain active motor exploration and mitigate the risk of learner dependency. While continuous force feedback effectively guides users along reference trajectories, it fundamentally alters the dynamics of the task. This physical intervention often induces the “guidance hypothesis” [13], where learners rely on robotic assistance and fail to internalize the task dynamics, resulting in poor long-term retention [14, williams2014motor]. Conversely, discrete haptic feedback acts as an informational cue without the physical constraints. Prior research demonstrates that vibrotactile feedback effectively conveys proprioceptive information under cognitive load [21], drives targeted biomechanical adaptations in dynamic gait retraining [1], and prompts voluntary limb movement in stroke rehabilitation [18]. By relying on discrete nudges, our framework requires learners to actively decode the haptic stimulus and execute an intrinsic motor plan, ensuring independent exploration of the redundant task manifold.
@@
Central to our approach is the recognition that motor learning is a latent process, in which observable performance is a noisy emission of the underlying, evolving motor learning and control strategy [20]. Drawing on human skill modeling literature [23, 20, 26], we formalize motor learning as a progression through discrete latent states [25], paralleling the cognitive, associative, and autonomous phases proposed by Fitts and Posner [2]. While standard Hidden Markov Models (HMM) have successfully tracked such transitions in skill acquisition [20, megali2006modelling, french2024determining, 3] and neural motor planning [8, 9], effective robotic nudge feedback requires understanding how external interventions influence these transitions. Thus, we employ a probabilistic predictive model of human motor learning using an Input-Output Hidden Markov Model (IOHMM). This structure allows us to causally link specific haptic nudges (inputs) to probabilities of skill improvement, effectively separating execution noise from true latent capability.
@@
Several mechanistic models have been proposed in the literature to explain the stochastic human motor learning process [19, 16, 7]. One salient feature of these models is the Markovian evolution of latent states as a function of certain inputs to the model. These models can be abstracted into a probabilistic finite-state machine defined over a quantized state space with event-triggered temporal sampling. The IOHMM operates on this abstraction and incorporates the resulting finite-state machine as its latent dynamics. This probabilistic framework allows us to treat the haptic tutoring as a sequential decision-making problem under uncertainty. Leveraging the IOHMM, we formulate a Partially Observable Markov Decision Process (POMDP) to derive an optimal nudging policy. Similar to prior work that utilized POMDPs to infer latent user fatigue during rehabilitation [6], our formulation models the evolution of latent motor skill over the duration of the motor task being performed by the learner. The objective is to prescribe a sequence of haptic feedback nudges that minimizes long-term performance cost, and implicitly guides the learner toward effective regions of the latent skill space. This enables a principled, data-driven approach, going beyond performance or expert-designed heuristics, for personalizing haptic feedback design in novel skill acquisition paradigms.
@@
Building on the preliminary findings of our POMDP-based optimal nudge feedback [kamboj2026toward], this work investigates the potential of the skill-informed haptic feedback design in accelerating skill acquisition and task performance for high-dimensional, de-novo motor tasks. We distinguish this study from previous work [kamboj2026toward] through the following advancements: (i) we introduce a new participant group training on a heuristic nudging policy to benchmark performance against no-feedback and POMDP-based optimal nudge feedback, and increase the sample size to better delineate the group effects, (ii) we employ additional analysis to evaluate how effectively different group participants are discovering the task space manifold, (iii) we analyzed the evolution of participants’ estimated latent skill states over the experiment trials, (iv) we analyze the participants’ motor output performance using a reaching error metric across the experiment trials, and (v) we expand upon the IOHMM employed to capture the human motor learning behavior, the training algorithm used, and provide interpretability to the estimated model parameters. In summary, our contributions are threefold as follows.
@@
    We introduce an IOHMM framework that captures the dynamics of high-dimensional motor learning with motor skill as a latent state, haptic nudge, and task type as inputs, and performance measures as noisy outputs.
@@
    We validate our approach through a human participant study, demonstrating that our POMDP-derived policy significantly expedites motor skill learning and enhances both endpoint accuracy and movement efficiency compared to both no-feedback and a heuristic feedback strategy.
@@
Figure 1: Target Capture Task: (a) shows a participant performing our target capture task with the SenseGlove DK1 exoskeleton strapped to their right hand. (b) and (c) show the cursor trajectories during the first and eighth blocks of the game with the red dots representing the target locations. The trajectories get straighter as the participant learns to control the cursor by the end of the experiment session. (d) shows how the performance metrics, Reaching Error and Straightness of Trajectory, are computed.
@@ 2 Background
This section details the experimental framework and the performance metrics used to evaluate human performance during the acquisition of a novel, high-dimensional motor skill. Subsequently, we outline the stochastic modeling approach adopted in this paper.
@@ 2.1 Experimental Setup
We adopt a Body-Machine Interface (BoMI) paradigm similar to those described in [15, 17]. Participants wore a SenseGlove DK1 (manufactured by SenseGlove, a Netherlands-based company), a non-invasive hand exoskeleton capable of tracking $20$ distinct degrees of freedom (DoF) across the hand joints. Additionally, the device can deliver independent vibrotactile haptic feedback to each of the five fingertips. This high-dimensional joint space is then projected into a lower-dimensional space, represented by a cursor on a screen. This projection is governed by the linear mapping
@@
where $\dot{\boldsymbol{x}}\in\mathbb{R}^{n}$ represents the velocity of the screen cursor, $\boldsymbol{u}=\dot{\boldsymbol{q}}\in\mathbb{R}^{m}$ denotes the finger joint velocities derived from the joint angles $\boldsymbol{q}$, and $C\in\mathbb{R}^{n\times m}$ serves as the BoMI mapping matrix. Participants are instructed to perform a target capture task [^1] A video footage of the target capture task is available at https://youtu.be/WIDRwqpE9Sk (see Fig. 1), in which they must coordinate their finger joint movements to navigate the cursor toward sequentially appearing targets. Because the mapping from the $20$-dimensional hand space to the $2$-dimensional cursor space is initially unknown to the user, the task enforces a high-dimensional learning process. Success requires the participant to discover and internalize specific coordinated hand movements consistent with the mapping matrix $C$. A haptic nudge, if prescribed to the participant, consists of a tactile stimulus: two sequential vibration bursts, each lasting 150150~ms, separated by 20002000~ms, delivered to the fingertip prescribed by the active feedback policy. The strength of this nudge is chosen such that it is clearly perceived by the participant, but doesn’t impact the finger motion. The experimental protocol is divided into three distinct phases:
  • Calibration: To customize the mapping, participants mimic a series of American Sign Language (ASL) alphabet hand postures. We apply Principal Component Analysis (PCA) to the recorded hand posture data and construct the mapping matrix $C$ using the first and the third principal components. The coordinate system is centered on the task space, represented as a $5\times 5$ unit game window, such that the user’s mean hand posture corresponds to the center of the game screen. One unit is the width of a grid square, computed to accommodate $1$-standard deviation of the calibration data points and to ensure all potential targets are comfortably within the user’s reach.
  • Familiarization: A brief acclimation period of approximately $6$ seconds is provided. During this time, participants are free to move the cursor within the task space (Fig. 1). This duration is intentionally short to allow users to get a feel for the system dynamics without inducing significant motor learning or adaptation effects.
  • Training: The primary gameplay phase comprises $8$ blocks, each containing $60$ individual trials. A trial consists of a single reaching movement toward a prescribed target; upon successful capture, a new target is immediately presented. A capture is registered if the cursor remains stable within the target’s grid square—defined as position changes of less than $0.0025$ units over $15$ consecutive samples at $100$ Hz. The targets are prescribed uniformly randomly from a set of $4$ targets $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3},\mathcal{T}_{4}\}$ at $\{(0.5,4.5),(2.5,2.5),(2.5,0.5),(4.5,4.5)\})$ units of the task space grid, respectively.
@@ 2.2 Performance Metrics
To comprehensively evaluate the motor learning process, we utilize endpoint accuracy, which measures task success, and movement efficiency, which characterizes the optimality of the underlying motor control strategy. In highly redundant tasks, learners frequently achieve high accuracy using suboptimal, compensatory movements. Therefore, evaluating both metrics is essential to verify true skill acquisition. Consequently, we quantify task performance using two primary metrics [7]: Reaching Error (RE) to evaluate endpoint accuracy, and Straightness of Trajectory (SoT) to evaluate movement efficiency (Fig. 1(d)).
  • RE is defined as the Euclidean distance between the cursor and the target center at the end of a trial. The end of the trial is defined as the instant when the cursor position did not change by more than $0.0025$ units for $15$ consecutive samples, or $2$ seconds after the start of the movement, whichever occurs first.
@@ 2.3 IOHMM and POMDP Framework
To model the stochastic dynamics of the user’s internal skill-state, we employ an Input-Output Hidden Markov Model (IOHMM) [bengio1996input]. This probabilistic graphical model describes the evolution of a discrete latent state sequence $\{s_{k}\in\mathcal{S}\}_{k\in\mathbb{N}}$, where $\mathcal{S}$ is the state space, and the transition to a future state $s_{k+1}$ depends on both the current state $s_{k}$ and an exogenous control input or action $a_{k}\in\mathcal{A}$, where $\mathcal{A}$ is the action space. Crucially, the true system state is inaccessible for direct measurement; it is hidden. Instead, inference is performed via observations $\{o_{k}\in\Omega\}_{k\in\mathbb{N}}$, where $\Omega$ is the observation space. These are modeled through the conditional probability distribution $\mathbb{P}(o|s,a)$, representing the likelihood of observing output $o$ given the system is in state $s$ and receives input $a$. We assume $\mathcal{S}$ and $\mathcal{A}$ are finite sets.
Given a historical dataset of input-output pairs $\{a_{k},o_{k}\}_{k\in\mathbb{N}}$, the IOHMM is trained to learn the state transition dynamics $T(s,a,s^{\prime})=\mathbb{P}(s^{\prime}|s,a)$, the observation probability matrix $O(s,a,o)=\mathbb{P}(o|s,a)$, and the initial state distribution. This parameter estimation is typically achieved via the extended Baum-Welch algorithm [bengio1996input], which adapts the standard Expectation-Maximization (EM) algorithm to account for the input-dependent structure of the IOHMM.
To transition from system modeling to active control, we extend the IOHMM into a Partially Observable Markov Decision Process (POMDP). The POMDP framework introduces an agent that optimizes the selection of inputs based on a defined objective. Formally, the POMDP is characterized by the tuple $\langle\mathcal{S},\mathcal{A},T,R,\Omega,O\rangle$. In this formulation, $\mathcal{S},\mathcal{A},T,\Omega,$ and $O$ retain their definitions from the IOHMM, while $R(s,a)$ represents the reward function. This function quantifies the agent’s performance goals, assigning positive scalar values to beneficial state-action pairs and penalties (costs) to undesirable outcomes, thereby guiding the optimal control policy.
The IOHMM-POMDP framework offers specific analytical and computational advantages for modeling and influencing the human motor learning process. Unlike deep reinforcement learning or recurrent neural networks for sequence modeling and control, IOHMM enables sample-efficient parameter estimation from limited human data while preserving interpretability. The state transitions and emission models allow us to explicitly quantify the effects of targeted haptic interventions (Section 5). Furthermore, unlike standard Markov Decision Processes (MDPs) that assume fully observable states, the POMDP formally accounts for execution noise. By decoupling noisy task performance metrics from discrete latent skill states, the derived policy optimizes for long-term motor adaptation rather than over-correcting for transient kinematic errors.